5x-2y=30
-2x+2y=0

Answer by mananth(16946)

5 x -2 y = 30 .............1

-2 x + 2 y = 0 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by 1
5 x -2 y = 30
-2 x + 2 y = 0
Add the two equations
3 x = 30
/ 3
x = 10
plug value of x in (1)
5 x -2 y = 30
50 -2 y = 30
-2 y = 30 -50
-2 y = -20
y = 10
x=10, y=10